A metal ball of mass 100g is released from rest at the top of a vertical cliff. It hits the ground after falling for 4 seconds and penetrates 10cm into the ground. Taking g=10m/s², determine the height of the cliff.
1
Expert's answer
2019-09-09T11:02:00-0400

To solve this problem, it is enough for us to know the time during which the thrown ball reached the ground.


"h = gt^2\/2""h = 10*4^2\/2 = 80m"